A box contain 10 blue marbles, 8 red marbles, 6 yellow marbles, and 5 white marbles. What is the probability that if you choose a marble from the box without looking or at random, you will choose a red marble? Show your solution.
a. 5/29
b. 6/29
c. 8/29
d. 10/29
1
Expert's answer
2021-08-16T16:52:20-0400

"P(red)=\frac{8}{10+8+6+5}=\frac{8}{29}."

Answer: c. 8/29